z = -1 is what percentile

I'm having trouble how to start on this can you explain it and solve it please
Answer #1

Answer

We can use z value to area table to find the corresponding percentile

We will be using the following table to convert z score to area, then we will multiply the area by 100, to get the percentile

2 187 505 32287 6880113 0026500943 2 to 0000 135689 999 0 5 0 5 0 5 0 5 0 5 0 5 0 32211000-1223

Now, we have to find the value of z = -1 in the above table

we get area = 0.1587 corresponding to z = -1

now multiplying the area by 100 to get the required percentile

we get

percentile = area*100 = 0.1587*100 = 15.87%

So, required percentile is 15.87%